OpenStudy (dannyrod2000):

2x^2+3x=(3x-1)^2-x^2 put in standard form

OpenStudy (anonymous):

I think you can do it. (a+b)^2 = a^2 +2ab + b^2

OpenStudy (dannyrod2000):

can you show me in steps

OpenStudy (anonymous):

2x^2 + 3x = (3x)^2 -(2)(1)(3x) + 1 - x^2. I just expaded the ones in the brackets. I think you can take care of the rest.

OpenStudy (dannyrod2000):

thanks

OpenStudy (anonymous):

:)

OpenStudy (dannyrod2000):

do i combine like terms?

OpenStudy (anonymous):

yup thats correct, the ones with the same power (degree) are considered to be like terms.

OpenStudy (dannyrod2000):

ok so i just combine and switch from one side to another?